OpenStudy (anonymous):

The unemployment rates for March in 2 similar sized towns are being compared. Town 1: of 1508 employed, 405 were laid off Town 2: of 1252 employed, 281 were laid off A) Test if there is any difference in the lay off rates at 5% level. . Please help me decide the order of steps. EOV first then Average????????? Ho:___________ .05 (2-tail) Ha:___________ Reject if: pvalue<.05 conclusion: b)is there diff in lay off rates, how large is difference??

OpenStudy (anonymous):

town1 unemployed = \[{405\over1058}\times100=?\] town2 unemployed = \[{281\over1252}\times100=?\]

OpenStudy (anonymous):

26.86 22.44

OpenStudy (anonymous):

I didn't multiply by 100 though thats new

OpenStudy (anonymous):

what is EOV now??

OpenStudy (anonymous):

when you multiply by 100, the result is in percentage

OpenStudy (anonymous):

equality of variance

OpenStudy (anonymous):

% ok yes right

OpenStudy (anonymous):

your calculations seem erroneous

OpenStudy (anonymous):

because its 1508, and you did 1058 for town 1 :)

OpenStudy (anonymous):

ooh :P

OpenStudy (anonymous):

but what I do not understand is how to use the F-test when there is just one information given.

OpenStudy (anonymous):

\[H_0:\sigma_1^2=\sigma_2^2\]

OpenStudy (anonymous):

can we find the \[\alpha\]

OpenStudy (anonymous):

ok i see about the F test is rejected then.

OpenStudy (anonymous):

yes it is .05

OpenStudy (anonymous):

α=.05

OpenStudy (anonymous):

this is exactly 5%

OpenStudy (anonymous):

yes that is alpha.

OpenStudy (anonymous):

you get \[\alpha=0.05\] for both?

OpenStudy (anonymous):

in our class we only use .05 for alpha α.

OpenStudy (anonymous):

but here we are required to find the alpha

OpenStudy (anonymous):

?

OpenStudy (anonymous):

alpha is always .05, i'm just unsure which program on the calculator ti it is?

OpenStudy (anonymous):

use the T-tables to find the "t" value for equal variances

OpenStudy (anonymous):

ok so this is equal. I was thinking it was UNequal?

OpenStudy (anonymous):



OpenStudy (anonymous):

we are using Null hypothesis what is sigma for H0 and what is sigma for Ha

OpenStudy (anonymous):

so do i use Z then? aghhh im lost these words al lmix together

OpenStudy (anonymous):

0 and 1

OpenStudy (anonymous):

we are trying to compare the "p" values with 0.05

OpenStudy (anonymous):

so is lay off rates the average.. ?

OpenStudy (anonymous):

ok logically speaking, \[p_1={405\over1508}=0.2686\\ p_2={281\over1252}=0.2244\\ p=|p_1-p_2|=0.0442<0.05 \] so, we reject the null hypothesis and the difference between the layoff rates = 0.0442=4.42%